# A Minimalist Optimizer Design for LLM Pretraining

### v1

@@ Abstract
Training large language models (LLMs) typically relies on adaptive optimizers such as Adam, which require significant memory to maintain first- and second-moment matrices, which are known as optimizer states. While recent works such as GaLore, Fira and APOLLO have proposed state-compressed variants to reduce memory consumption, a fundamental question remains: What is the minimal amount of optimizer state that is truly necessary to retain state-of-the-art performance in LLM pretraining? In this work, we systematically investigate this question using a bottom-up approach. We find that two (memory- and compute-efficient) optimization techniques are particularly effective: (1) column-wise gradient normalization significantly boosts the performance of plain SGD without requiring momentum; and (2) adding first-order momentum only to the output layer – where gradient variance is highest – yields performance competitive with fully adaptive methods such as Muon. Based on these insights, we propose SCALE (Stochastic Column-normAlized Last-layer momEntum), a new optimizer that combines column-normalized SGD with last-layer momentum, where column normalization refers to normalizing the gradient along the output dimension. Across multiple LLaMA models (60M–1B), SCALE matches or exceeds the performance of Adam while using only 35–45% of the total memory. It also consistently outperforms memory-efficient optimizers such as GaLore, Fira and APOLLO, making it a strong candidate for large-scale pretraining under memory constraints. For LLaMA 7B model, SCALE outperforms the state-of-the-art method APOLLO, in terms of both perplexity and memory consumption. In addition, our method serves as a minimalist baseline for a more sophisticated optimizer design. Code is available at this link.
@@ 1 Introduction
Adaptive optimizers such as RMSProp (Hinton et al., 2012), Adam (Kingma and Ba, 2015) are the default optimizers for large-scale attention-based deep neural networks, particularly large language models (LLMs). While effective, these optimizers incur significant memory overhead due to their reliance on maintaining both first- and second-moment estimates of the gradient, which are also known as optimizer states. For example, Adam requires storing two additional tensors per parameter tensor, tripling the memory usage compared to vanilla stochastic gradient descent (SGD).
On the other hand, despite its superior memory efficiency, the vanilla SGD performs poorly when applied directly in LLM training, due to the absence of adaptive scaling in the update step (See Figure 2 for experiment results, also see Zhao et al. (2025); Zhang et al. (2020b)). This has motivated a wave of recent research focused on developing memory-efficient alternatives to Adam that aim to retain its performance while reducing memory consumption. These include compression-based algorithms such as GaLore (Zhao et al., 2024) and Adam-Mini (Zhang et al., 2024), as well as novel designs such as Muon (Jordan et al., 2024), Sophia (Liu et al., 2024), SGD-SaI (Xu et al., 2024), Scion Pethick et al. (2025) and SWAN (Ma et al., 2024). These approaches often introduce new algorithmic components—such as different forms of gradient normalization, momentum variants, or second-order approximations—and combine them in various ways.
However, despite the growing literature on memory-efficient optimizers, there has been no systematic study to identify which specific algorithmic components or techniques are most essential for designing highly-performed yet minimal-memory optimizers. For instance, are both first- and second-order momentum terms strictly necessary for effective training? Among the many forms of gradient normalization, which subset of them strike the best trade-off between performance, memory, and computational cost? In the absence of such a principled investigation, it remains unclear how to balance optimizer complexity against memory savings.
@@
What is the minimum subset of optimizer states beyond the SGD that is required to achieve state-of-the-art pretraining performance?
In this work, we pursue this question through a bottom-up, minimalist approach. Rather than combining a large number of techniques, we systematically evaluate the contributions of basic components – such as normalization and momentum – to identify the minimal set of techniques (on top of vanilla SGD) that are necessary to match Adam-level performance. Concretely, we:
(1) Identify two key components that can effectively enhance the performance of vanilla SGD: gradient normalization and first-order momentum;
(2) Analyze how these components can be adapted in the most memory- and compute-efficient manners;
(3) Justify our design choices using a combination of theoretical insights and empirical evidence.
Our study suggests that two techniques, when used together, are particularly effective: (i) column-wise gradient normalization, which improves SGD performance significantly without requiring momentum; and (ii) adding first-order momentum exclusively to the output layer, where gradient variance is highest (see Figure 3). These insights lead us to propose Stochastic Column-normalized Last-layer momentum (SCALE), a minimalist optimizer that combines these two techniques, which requires roughly the same amount of memory as compared to vanilla SGD. For example, for 1B (resp. 7B) model, SCALE only requires 10% (resp. 2%) more memory as compared to vanilla SGD. Meanwhile, SCALE achieves competitive performance as compared with SOTA optimizers Adam and Muon, with only 35% and 52% of the memory cost for training 1B models, respectively. See Figure 1 for an illustration of performance and memory trade-off among different SOTA algorithms.
@@ 1.1 Related Works
Memory-efficient variations of Adam. A recent line of works aims to improve memory efficiency of Adam, compressing the historical states stored, namely the first- and second-order statistics by the use of gradient projections. GaLore (Zhao et al., 2024), being one of the pioneering works, stores the states in a low-rank subspaces that captures most gradient information. Fira (Chen et al., 2024) achieves superior performance than GaLore by re-introducing full-rank information to the low-rank gradients. APOLLO (Zhu et al., 2024) constructs the update based on gradient scaling factors that are estimated from the ratio between the low-dimensional gradient and the low-dimensional Adam update; APOLLO-Mini is a memory efficient version of APOLLO by estimating in a rank-1 subspace. GRASS (Muhamed et al., 2024) improves GaLore by designing sparse projection matrices guided by the norm of the rows of the gradients. SlimAdam (Kalra et al., 2025) compresses second-order moments based on signal-to-noise analysis. Some other methods group parameters into blocks and apply block-wise updates (Luo et al., 2024; Ramesh et al., 2024; Pan et al., 2024), block-wise scaling (Zhang et al., 2024), channel-wise scaling (Zhu et al., 2025), to further reduce the memory costs.
Towards stateless optimizers. More recent works start to question the necessity of optimizer states as required by Adam. Adafactor (Shazeer and Stern, 2018) is perhaps the first work to record the momentum of the column and row norms of the gradient matrix, reducing the memory of optimizer states from $\mathcal{O}(d^{2})$ to $\mathcal{O}(d)$ (assuming that the gradient matrix is $d$ by $d$). SWAN (Ma et al., 2024) combines two gradient normalizations (singular-value and row-wise normalizations, see Section 2), which matches the performance of Adam (when applied only to the intermediate layers while the first and last layers still use Adam. See Section 4 for details). Zhao et al. (2025) demonstrates SGD with signed momentum is able to recover the performance of Adam. SGD-SaI (Xu et al., 2024) verifies that proper learning‑rate scaling at initialization is sufficient to achieve good performance.
Novel optimizers with better performance than Adam. Another line of work focuses on proposing new optimizers to achieve improved convergence properties compared to Adam. SPAM and Stable‑SPAM (Huang et al., 2025b, a) reset momentum and clip spike gradients, stabilizing mixed‑precision LLM training and yield better performance than Adam. Sophia (Liu et al., 2024) accelerates training by leveraging an estimated diagonal Hessian. Muon (Jordan et al., 2024) proposes gradient orthogonalization to further stabilize and enhance LLM training. Other approaches (Bernstein and Newhouse, 2024; Pethick et al., 2025; Kalra et al., 2025) explore various (layer-wise) normalization schemes to design better adaptive optimizers.
@@ 2 Methodology
We begin by reviewing several key techniques used in popular optimizers such as Adam and Muon, which have been shown to accelerate pretraining performance for LLMs. This understanding will be critical for our subsequent minimalist design for a memory-efficient optimizer.
Denote the optimization problem of the LLM pretraining as
@@
where $\ell$ is the loss function, usually taken as the cross entropy loss of predicting the next token, $\theta=[\theta_{1},...,\theta_{L}]$ is the model trainable parameters, with $\theta_{l}$ the $l$-th layer, $l=1,2,...,L$. With attention-based network, we can simply assume that each $\theta_{l}\in\mathbb{R}^{d_{l,\text{in}}\times d_{l,\text{out}}}$ is a weight matrix, with the input dimension $d_{l,\text{in}}$ and output dimension $d_{l,\text{out}}$. Here $\xi_{i}$ with $i=1,...,n$ represents training data samples and $n$ is the training data size.
@@
where $B$ is batch size and $\eta_{t}$ is the learning rate. Although memory-efficient, plain SGD performs poorly in LLM training due to the lack of adaptive scaling  (Zhao et al., 2025), and we verify this in Figure 2 where we run SGD and Adam on LLaMA 130M pretraining task (see Section 4 for the details of our experiment settings). Therefore we will not include the perplexity result for SGD in the subsequent experiments.
@@
Figure 2: Comparison of SGD and Adam training loss and evaluation perplexity curves on LLaMA 130M model. Clearly, SGD is not converging to any reasonable level of perplexity. The Adam and SGD learning rates are 3e-3 and 0.1, respectively. We search with multiple learning rates for SGD, for lower ones the loss decreases even slower and higher ones would cause the training to diverge.
In contrast, adaptive algorithms such as Adam (Kingma and Ba, 2015) update parameters using a more sophisticated scheme (where $\odot$ represents element-wise product):
@@
Numerous theoretical works such as Reddi et al. (2019); Zhang et al. (2022b, a) attempted to explain why Adam outperforms SGD for transformers, though consensus remains elusive. A straightforward decomposition of Adam identifies two essential components.
@@
    $\theta^{t+1}=\theta^{t}-\eta_{t}\frac{g^{t}}{\sqrt{g^{t}\odot g^{t}}}=\theta^{t}-\eta_{t}\mathop{\mathrm{missing}}{sign}(g^{t})$ (2.4)
    where $\mathop{\mathrm{missing}}{sign}$ stands for taking the sign for each element.
@@
This leads to the central question: How much of Adam’s benefit arises from gradient normalization versus EMA? We note that gradient normalization does not require maintaining state, whereas EMA does. Therefore, in the next sections, we will examine each of these components separately and discuss their effectiveness. We will start from the gradient normalization, as the EMA step will introduce extra memory no matter how hard we compress it. Then we will discuss how to use EMA in a more memory-friendly way to further boost the performance of the optimizer.
@@ 2.1 Gradient normalization
Gradient normalization is a critical component for large-scale, efficient pretraining, and has remained indispensable even in recent (near)-stateless optimizers (Ma et al., 2024; Huang et al., 2025a; Zhu et al., 2024). It is known to accelerate escape from saddle points (Levy, 2016; Murray et al., 2019), improve the effective smoothness of the objective (when the Hessian norm is upper bounded in terms of the gradient norm) (Zhang et al., 2020a; Kunstner et al., 2023), stabilize gradient distributions (Ma et al., 2024), and provide robustness to heavy-tailed noise (Sun et al., 2024).
In this section, we build upon the framework presented in Bernstein and Newhouse (2024); Pethick et al. (2025) to explore a family of gradient normalization methods within the context of trust-region steepest descent. Denote $G^{t}$ the stochastic gradient of current iteration (we switch to upper letters since we assume the weight blocks $\theta_{l}$ in (2.1) are matrices). It has been observed that different normalization techniques arise from using different matrix norms in the following steepest descent formulation:
$\Theta^{t+1}=\Theta^{t}+\eta_{t}\Delta^{t},\quad\text{ where }\Delta^{t}=\mathop{\mathrm{missing}}{argmin}_{\|\Delta\|\leq\rho}\ \langle G^{t},\Delta\rangle$ (2.6)
where $G^{t}\in\mathbb{R}^{m\times n}$ denotes the (stochastic) gradient at iteration $t$. In particular, we use $\|A\|_{p\rightarrow q}$ to represent the operator norm of matrix $A$ induced by vector norm $\ell_{p}$ and $\ell_{q}$, i.e., $\|A\|_{p\rightarrow q}:=\max_{\|x\|_{p}\leq 1}\|Ax\|_{q}$. We can derive the following results based on Bernstein and Newhouse (2024) [^1] It is established in Bernstein and Newhouse (2024) that $\mathop{\mathrm{missing}}{argmin}_{\|\Delta\|\leq\rho}\ \langle G,\Delta\rangle=\frac{\rho\nu}{\|G\|_{*}}\mathop{\mathrm{missing}}{argmin}_{x}\ \langle G,\Delta\rangle+\frac{\nu}{2}\|\Delta\|^{2}$ which indicates that we can also put all the normalization in the context of quadratic minimization.:
$\mathop{\mathrm{missing}}{argmin}_{\|\Delta\|_{2\rightarrow 2}\leq 1}\ \langle G,\Delta\rangle=-UV^{\top},\text{ where }G=U\Sigma V^{\top}\text{ is the SVD},$ (2.7)
$\mathop{\mathrm{missing}}{argmin}_{\|\Delta\|_{1\rightarrow 2}\leq 1}\ \langle G,\Delta\rangle=-\left[\frac{\text{col}_{1}(G)}{\|\text{col}_{1}(G)\|},...,\frac{\text{col}_{n}(G)}{\|\text{col}_{n}(G)\|}\right],$
$\mathop{\mathrm{missing}}{argmin}_{\|\Delta\|_{2\rightarrow\infty}\leq 1}\ \langle G,\Delta\rangle=-\begin{bmatrix}\frac{\text{row}_{1}(G)}{\|\text{row}_{1}(G)\|}\\ \cdots\\ \frac{\text{row}_{m}(G)}{\|\text{row}_{m}(G)\|}\end{bmatrix},$
$\mathop{\mathrm{missing}}{argmin}_{\|\Delta\|_{1\rightarrow\infty}\leq 1}\ \langle G,\Delta\rangle=-\mathop{\mathrm{missing}}{sign}(G),$
which we refer to as singular-value, column-wise, row-wise and sign normalization, respectively. Here we assume that $m$ is the input dimension and $n$ is the output dimension, therefore row- and column-normalizations correspond to normalizing along the input and output dimensions, respectively.
It is worth mentioning that multiple existing works can be summarized in this gradient normalization framework. For example, Sign-SGD/Adam utilize $\|\cdot\|_{1\rightarrow\infty}$ (sign normalization) to achieve gradient normalization, Muon (Jordan et al., 2024) utilizes $\|\cdot\|_{2\rightarrow 2}$ norm (singular-value normalization), whereas SCION (Pethick et al., 2025) and SlimAdam (Kalra et al., 2025) apply different norms for different layers to achieve better performance.
Computational cost of different normalizations. In terms of the computational cost, the normalization techniques discussed in (2.7) can be quite different. In particular, singular-value normalization requires computing full SVD, which is the most time consuming compared to the rest. Even though efficient approximation methods of SVD has been studied in Jordan et al. (2024); Ma et al. (2024) (e.g., the Newton-Schulz (NS) procedure), they are still much more time consuming as compare with the other three normalization techniques, see Table 4 for a test on the time required for different normalizations [^2] In all following experiments, we conduct singular-value normalization with the Newton-Schulz (NS) procedure as in Jordan et al. (2024); Ma et al. (2024)..
@@
Table 1: Time (ms) consumed by each of the normalization methods on a torch matrix tensor with dimension $d_{\text{in}}=d_{\text{out}}=d$, testing on a single NVIDIA A40 GPU. Here the singular-value normalization are computed both exactly (first row, using torch.linalg.svd directly) and inexactly using Newton-Schulz (NS) iteration (second row, see Jordan et al. (2024) for details) [^4] The difference of column- and row-wise normalization may originate from the way PyTorch stores tensors with strides, see this link..
Experimental comparison of SGD with different gradient normalizations. We conduct the preliminary experiment of pretraining LLaMA models (See Section 4 for the experiment setting) using SGD (2.6) with different normalizations applied to all the layers as specified in (2.7). We report the pretraining perplexities in Table 2, and we notice that all the normalizations improve over SGD, however none of them alone could match the performance of Stable-SPAM (which is a stabilized version of Adam) or Muon, two of the state-of-the-art algorithms. In particular, the four normalization methods can be categorized into two groups based on their performance: singular-value and column-wise normalization achieve better results than row-wise and sign normalization. We thus proceed with the group of normalizations with better performance, namely the singular-value and column-wise normalizations.
@@
Stable-SPAM 28.77 22.20 16.80
Muon 28.86 22.20 16.70
singular-value 34.15 25.25 18.73
@@
Table 2: Preliminary experiments results (perplexity) of pretraining LLaMA models on C4 dataset, using different norms for steepest gradient descent (2.6) methods as specified in (2.7). For the singular-value normalization, we use the inexact Newton-Schulz (NS) iteration (again see Jordan et al. (2024) for details) for fast approximation.

### v2

@@ Abstract
Training large language models (LLMs) typically relies on adaptive optimizers such as Adam, which introduce extra operations and require significant more memory to maintain first- and second-order moments than SGD. While recent works such as GaLore, Fira and APOLLO have proposed state-compressed variants to reduce memory consumption, a fundamental question remains: What are the minimum modifications to plain SGD needed to match state-of-the-art pretraining performance? We systematically investigate this question using a bottom-up approach, and identify two simple yet highly (memory- and compute-) efficient techniques: (1) column-wise gradient normalization (normalizing the gradient along the output dimension), which boosts SGD performance without momentum; and (2) applying first-order momentum only to the output layer, where gradient variance is highest. Combining these two techniques lead to SCALE (Stochastic Column-normAlized Last-layer momEntum), a simple optimizer for memory efficient pretraining. Across multiple LLaMA models (60M–1B), SCALE matches or exceeds the performance of Adam while using only 35–45% of the total memory. It also consistently outperforms memory-efficient optimizers such as GaLore, Fira and APOLLO, making it a strong candidate for large-scale pretraining under memory constraints. For LLaMA 7B model, SCALE outperforms the state-of-the-art memory-efficient methods APOLLO and Muon, in terms of both perplexity and memory consumption. Code is available at this link.
@@ 1 Introduction
Adaptive optimizers such as RMSProp (hinton2012neural), Adam (kingma2015adam) are the default optimizers for large-scale attention-based deep neural networks, particularly large language models (LLMs). While effective, these optimizers incur significant memory overhead due to their reliance on maintaining both first- and second-moment estimates of the gradient, which are also known as optimizer states. Specifically, Adam requires storing two additional gradient states per parameter, tripling the memory usage compared to vanilla stochastic gradient descent (SGD).
On the other hand, despite its superior memory efficiency, the vanilla SGD performs poorly when applied directly to LLM training, due to the absence of adaptive scaling in the update step (See Figure 2 for experiment results, also see zhao2025deconstructing; zhang2020adaptive). This has motivated a wave of recent research focused on developing memory-efficient alternatives to Adam that aim to retain its performance while reducing memory consumption. These include compression-based algorithms such as GaLore (zhao2024galore), Fira chen2024fira and APOLLO zhu2025apollo. Novel optimizers, such as Muon (jordan2024muon), Scion (pethick2025scion) and SWAN (ma2024swan), also by design require less memory than Adam. These approaches often introduce new algorithmic components—such as different forms of gradient normalization, whitening and rescaling, momentum variants—and combine them in various ways.
However, despite the growing literature on memory-efficient optimizers, there has been no systematic study to identify which specific algorithmic components or techniques are most essential for designing highly performant yet minimal-memory optimizers. For instance, are both first- and second-order momentum terms strictly necessary for effective training? Among the many forms of gradient normalization, which subset of them strike the best trade-off between performance, memory, and computational cost? In the absence of such a principled investigation, it remains unclear how to balance optimizer complexity against memory savings.
@@
Can we design a memory efficient optimizer with minimum modifications to plain SGD that achieves state-of-the-art pretraining performance?
In this work, we address this question through a bottom-up, minimalist approach. Rather than starting from existing adaptive optimizers, we systematically evaluate the role of fundamental components, namely normalization and momentum, to determine the smallest set of techniques needed to bridge the gap between vanilla SGD and Adam, with a memory efficiency focus. Towards this end, we perform extensive experiments to identify key components (such as different forms of normalization and different levels of momentum) that can effectively enhance the performance of vanilla SGD with minimum memory overhead. We then justify the design choices using a combination of theoretical insights and empirical evidence.
Contributions. Our study suggests that two techniques, when used together, are particularly effective: (i) Among various gradient normalization techniques in the literature, column-wise gradient normalization (normalizing along the output dimension) can significantly boost SGD performance, while having simple closed-form solution and with no extra memory required; and (ii) adding first-order momentum exclusively to the output layer, where gradient variance is highest (see Figure 4), is surprisingly effective to further boost training performance with minimum memory overhead. Combining these two techniques lead to Stochastic Column-normalized Last-layer momentum (SCALE), a memory efficient optimizer which requires roughly the same amount of memory as vanilla SGD. For example, for 1B (resp. 7B) model, SCALE only requires 10% (resp. 2%) more memory as compared to vanilla SGD. Meanwhile, SCALE achieves similar performance as compared with SOTA optimizers Adam and Muon, with only 35% and 52% of the memory cost for training 1B models, respectively. Comparing to other memory efficient optimizers, SCALE achieves superior performance to GaLore, Fira and Apollo, with only 59% of the memory cost for training 1B models. See Figure 1 for an illustration of performance and memory trade-off among different SOTA algorithms.
@@ 1.1 Related Works
Memory-efficient variations of Adam. A recent line of works aims to improve the memory efficiency of Adam by compressing the historical states stored, namely the first- and second-order statistics. Early works include Adafactor (shazeer2018adafactor), which estimates that optimizer states using per-row and per-column moving averages, SM3 (anil2019memory), a memory efficient AdaGrad via grouping second-order momentum, and CAME (luo2023came), which improves upon Adafactor via matrix factorization. A number of recent works use gradient projections to compress the optimizer states. GaLore (zhao2024galore), being one of the pioneering works, stores the states in a low-rank subspaces that captures most gradient information. Fira (chen2024fira) achieves superior performance than GaLore by re-introducing full-rank information to the low-rank gradients. APOLLO (zhu2025apollo) constructs the update based on gradient scaling factors that are estimated from the ratio between the low-dimensional gradient and the low-dimensional Adam update; APOLLO-Mini is a memory efficient version of APOLLO by estimating in a rank-1 subspace. GRASS (muhamed2024grass) improves GaLore by designing sparse projection matrices guided by the norm of the rows of the gradients. SlimAdam (kalra2025can) compresses second-order moments based on signal-to-noise analysis. Some other methods group parameters into blocks and apply block-wise updates (luo2024badam; ramesh2024blockllm; pan2024lisa) or block-wise scaling (zhang2024adammini) to further reduce the memory costs.
Towards removing optimizer states. More recent works start to directly remove the optimizer states as required by Adam. Muon (jordan2024muon; liu2025muon) has shown impressive results using only first order momentum and orthogonalizing it via an iterative algorithm. Scion (pethick2025scion) explores various (layer-wise) normalization schemes for better training performance. SWAN (ma2024swan) and its variant (scetbon2025gradient) combine two normalizations (see Section 2 for details) directly to the gradient, matching the performance of Adam (when applied only to the intermediate layers while the first and last layers still use Adam. See Section 4 for details). zhao2025deconstructing demonstrates SGD with signed momentum is able to give similar performance to Adam. SGD-SaI (xu2024no) verifies that proper learning‑rate scaling at initialization is sufficient to achieve good performance.
@@ 2 Methodology
We begin by reviewing several key techniques used in popular optimizers such as Adam and Muon, which have been shown to accelerate pretraining performance for LLMs. This understanding will be critical for our subsequent minimalist design for a memory-efficient optimizer. Denote the optimization problem of the LLM pretraining as
@@
where $\ell$ is the loss function, usually taken as the cross entropy loss of predicting the next token, $\theta=[\theta_{1},...,\theta_{L}]$ is the model trainable parameters, with $\theta_{l}$ the $l$-th layer, l=1,2,,Ll=1,2,...,L. With attention-based network, we can simply assume that each $\theta_{l}\in\mathbb{R}^{d_{l,\text{in}}\times d_{l,\text{out}}}$ is a weight matrix, with the input dimension $d_{l,\text{in}}$ and output dimension $d_{l,\text{out}}$. Here $\xi_{i}$ with i=1,,ni=1,...,n represents training data samples and $n$ is the training data size.
@@
where $B$ is batch size and $\eta_{t}$ is the learning rate. Although memory-efficient, plain SGD performs poorly in LLM training due to the lack of adaptive scaling  (zhao2025deconstructing), and we verify this in Figure 2 where we run SGD and Adam on LLaMA 130M pretraining task (see Section 4 for the details of our experiment settings). Therefore we will not include the perplexity result for SGD in the subsequent experiments.
@@
Figure 2: Comparison of SGD and Adam training loss and evaluation perplexity curves on LLaMA 130M model. Clearly, SGD is not converging to any reasonable level of perplexity. The Adam and SGD learning rates are 3e-3 and 0.1, respectively. We search with multiple learning rates for SGD, for lower ones the loss decreases even slower and higher ones cause the training to diverge.
In contrast, adaptive algorithms such as Adam (kingma2015adam) update parameters using a more sophisticated scheme (where $\odot$ represents element-wise product):
@@
Numerous theoretical works such as reddi2019convergence; zhang2022adamcanconverge; yushun2022doesadamconverge attempted to explain why Adam outperforms SGD for transformers, though consensus remains elusive. A straightforward decomposition of Adam identifies two essential components.
@@
    $\theta^{t+1}=\theta^{t}-\eta_{t}\frac{g^{t}}{\sqrt{g^{t}\odot g^{t}}}=\theta^{t}-\eta_{t}\mathop{\mathrm{sign}}(g^{t})$ (2.4)
    where $\mathop{\mathrm{sign}}$ stands for taking the sign for each element.
@@
Combining (2.4) and (2.5) results in the original Adam update (2.3).
This leads to the question: How much of Adam’s benefit arises from gradient normalization versus EMA? We note that gradient normalization does not require maintaining state, whereas EMA does. Therefore, in the next sections, we will examine each of these components separately and discuss their effectiveness. We will start from the gradient normalization, as the EMA step will introduce extra memory no matter how hard we compress it. Then we will discuss how to use EMA in a more memory-friendly way to further boost the performance of the optimizer.
@@ 2.1 Gradient normalization
Gradient normalization is a critical component for large-scale, efficient pretraining, and has remained indispensable even in recent (near)-stateless optimizers (ma2024swan; huang2025stablespam; zhu2025apollo). It is known to accelerate escape from saddle points (levy2016power; murray2019revisiting), improve the effective smoothness of the objective (when the Hessian norm is upper bounded in terms of the gradient norm) (Zhang2020Why; kunstner2023noise), stabilize gradient distributions (ma2024swan), and provide robustness to heavy-tailed noise (sun2024gradient).
Denote $G\in\mathbb{R}^{d_{\rm in}\times d_{\rm out}}$ the stochastic gradient (we switch to upper letters since we assume the weight blocks $\theta_{l}$ in (2.1) are matrices). There is a variety of gradient normalization schemes, including
$\begin{array}[]{rl}\text{Singular-value normalization:}&UV^{\top},\quad\text{ where }G=U\Sigma V^{\top}\ (\text{SVD})\\[6.0pt] \text{Column-wise normalization:}&\Big[\tfrac{\text{col}_{1}(G)}{\|\text{col}_{1}(G)\|},\dots,\tfrac{\text{col}_{n}(G)}{\|\text{col}_{n}(G)\|}\Big]\\[6.0pt] \text{Row-wise normalization:}&\Big[\tfrac{\text{row}_{1}(G)^{\top}}{\|\text{row}_{1}(G)\|},\dots,\tfrac{\text{row}_{m}(G)^{\top}}{\|\text{row}_{m}(G)\|}\Big]^{\top}\\[6.0pt] \text{Sign normalization:}&\operatorname{sign}(G)\end{array}$ (2.6)
Here we assume that $d_{\rm in}$ is the input dimension and $d_{\rm out}$ is the output dimension. Therefore row- and column-normalizations correspond to normalizing along the input and output dimensions, respectively. It can be verified that different normalization schemes arise naturally from the steepest ascent direction under different matrix norms (bernstein2024modular; bernstein2024old; pethick2025scion; bernstein2025modular). Many existing optimizers employ different normalization techniques. For example, Sign-SGD/Adam utilize $\|\cdot\|_{1\rightarrow\infty}$ (sign normalization) [^1] Here we use the matrix induced norm. For any matrix $A\in\mathbb{R}^{n\times m}$, define $\|A\|_{a\rightarrow b}=\max_{x\in\mathbb{R}^{m}}\frac{\|Ax\|_{b}}{\|x\|_{a}}$ where $\|\cdot\|_{a}$ and $\|\cdot\|_{b}$ are vector $\ell_{a}$ and $\ell_{b}$ norms. to achieve gradient normalization, Muon (jordan2024muon) utilizes $\|\cdot\|_{2\rightarrow 2}$ norm (singular-value normalization), whereas SCION (pethick2025scion) and SlimAdam (kalra2025can) apply different norms for different layers to achieve better performance. Below we systematically analyze the effect of each normalization scheme separately.
@@
Table 1: Time (ms) consumed by each of the normalization methods on a torch matrix tensor with dimension $d_{\text{in}}=d_{\text{out}}=d$, testing on a single NVIDIA A40 GPU. Here the singular-value normalization are computed both exactly (first row, using torch.linalg.svd directly) and inexactly using Newton-Schulz (NS) iteration (second row, see jordan2024muon for details) [^3] The difference of column- and row-wise normalization may originate from the way PyTorch stores tensors with strides, see this link..
@@
Adam (Stable-SPAM) 28.77 22.20 16.80
singular-value (NS) 34.15 25.25 18.73
@@
Table 2: Results (perplexity) of pretraining LLaMA models on C4 dataset, using different gradient normalizations, as specified in (2.6). For the singular-value normalization, we use the inexact Newton-Schulz (NS) iteration (again see jordan2024muon for details) for fast approximation.
Computational cost of different normalizations. In terms of the computational cost, the normalization techniques discussed in (2.6) can be quite different. In particular, singular-value normalization requires computing full SVD, which is the most time consuming compared to the rest. Even though efficient approximation methods of SVD has been studied in jordan2024muon; ma2024swan (e.g., the Newton-Schulz (NS) procedure), they are still much more time consuming as compare with the other three normalization techniques, see Table 3 for a test on the time required for different normalizations.
Experimental comparison of SGD with different gradient normalizations. We conduct the preliminary experiment of pretraining LLaMA models (See Section 4 for the experiment setting) using SGD with different normalizations applied to all the layers as specified in (2.6). We report the pretraining perplexities in Table 2, and we notice that all the normalizations improve over SGD, however none of them alone could match the performance of Stable-SPAM (which is a stabilized version of Adam). In particular, the four normalization methods can be categorized into two groups based on their performance: singular-value and column-wise normalization achieve better results than row-wise and sign normalization. We thus proceed with the group of normalizations with better performance, namely the singular-value and column-wise normalizations.
(a) Row-wise normalized LM-head gradients
(b) Column-wise normalized LM-head gradients
Figure 3: We present the histograms of the LM-head gradients after applying row-wise (a) and column-wise normalization (b). The gradients are from the 1000th training iteration of a LLaMA 130M model. It can be seen from figure (a) that row-wise results into some very high gradient values (up-to range 150 in absolute value) that we find to destabilize training.
Why row-wise normalization performs poorly. Table 2 shows that row-wise gradient normalization consistently achieves worse performance than column-wise. We identified that this is largely attributed to the effect of row-wise normalization to the last layer’s (LM-head) gradients. Figure 3 shows that after row-wise is applied to the LM-head, some gradients obtain very high absolute values (up-to 150 in this case), which destabilizes the training. In contrast, column-wise results into a gradient distribution without extreme values, aiding stable training. We provide further discussion about the importance of column-wise for the last layer in Appendix A.9.